b1az3:

Fill in the blank below. After the expression is simplified as much as possible, x is raised to what exponent?

b1az3:

AZ:

You need to use the following rule \(\Large (a^b)^c = a^{bc}\) so what is \(\Large (x^8)^{\frac{3}{4}} = \) basically, multiply the exponents what is 8 * 3/4

b1az3:

hold up

b1az3:

x^6

b1az3:

but it isn't it for some reason

snowflake0531:

it is

b1az3:

i think it's my school

snowflake0531:

?_? Explain~

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @b1az3 x^6 \(\color{#0cbb34}{\text{End of Quote}}\) that is indeed correct, Kyle finally some progress

b1az3:

i swear it will be wrong again

AZ:

If it's wrong, then it won't be you who's wrong but instead the teacher who is wrong

snowflake0531:

Ye With the rule #1 that Angle's always correct, AZ's also always correct lol